# Revenue Maximization Under Sequential Price Competition Via The Estimation Of $s$-Concave Demand Functions

### v4

@@ Appendix A Summary of individual information
We now construct an estimator of $\psi_{i,\bm{\theta}}$ and establish, in Theorem D.6, a uniform convergence result over any bounded subset $K\subset\mathcal{U}$ (where $\mathcal{U}$ is defined in (5)). Our approach draws on techniques from mosching2020monotone; dumbgen2004consistency, which study uniform convergence in the context of shape-constrained estimation. A key assumption in these works – one that does not require smoothness of the unknown nonparametric function – is that the density of the design points, $f_{w}$, is bounded away from zero on $\mathcal{U}$. This ensures that the design points $w^{(t)}_{i}\in\mathcal{U}$ are asymptotically dense within every subinterval of $\mathcal{U}$, which in turn is sufficient for establishing uniform convergence of the estimator on such intervals. In our setting, $f_{w}$ is supported on all of $\mathbb{R}$, and thus it is bounded away from zero on any bounded set $K$. Consequently, the uniform convergence result applies to any bounded subset $K\subset\mathcal{U}$.
Before establishing the convergence result, we must first characterize the shape constraints satisfied by the family of functions $\{\psi_{i,\bm{\theta}}(\cdot):\bm{\theta}\in\mathbb{S}_{N-1}\}$. In Proposition D.4 (with proof deferred to Appendix F), we derive a closed-form expression for $\psi_{i,\bm{\theta}}(\cdot)$ and show that both the monotonicity and the $s_{i}=c_{i}-1$-concavity of the original function $\psi_{i}$ are preserved by $\psi_{i,\bm{\theta}}$, uniformly over $\bm{\theta}\in\mathbb{S}_{N-1}$.
@@
Under 3.1 and 5.1 we have the following properties:
@@
    If also 3.4 holds (i.e. $\psi_{i}$ is $s_{i}$-concave for some $s_{i}>-1$ as it arises from Proposition 3.5) then $\psi_{i,\bm{\theta}}$ is $s_{i}$-concave for any $\bm{\theta}\in\mathbb{S}_{N-1}$.
By definition of $s$-concavity, from point (5) in Proposition D.4 we have that for every $\bm{\theta}\in\mathbb{S}_{N-1}$ there exists a concave function $\phi_{i,\bm{\theta}}$ that $\psi_{i,\bm{\theta}}=h_{s_{i}}\circ\phi_{i,\bm{\theta}}$, where $h_{s_{i}}$ is defined in (4), and $s_{i}=c_{i}-1$ is known by 3.4. We could then estimate $\phi_{i,\bm{\theta}}(u)$ using LSE under concavity restriction, that is
@@
where we recall $w^{(t)}_{i}\triangleq\langle\bm{\theta},\mathbf{p}^{(t)}\rangle$. The interval $\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}]$ is such that $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}]=h_{s_{i}}([\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}])$. Note that $\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}]$ is known because $\underline{B}_{\psi_{i}}$ and $\bar{B}_{\psi_{i}}$ are known by 3.1 (recll that, by point (2) in Proposition D.4, the bounds $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}$ are inherited from $\psi_{i}$ to $\psi_{i,\bm{\theta}}$). We want to highlight that for any fixed $\bm{\theta}\in\mathbb{S}_{N-1}$, the problem (19) is not infinite dimensional, indeed we only need to recover $n$ variables $(\phi_{i}(w_{1}^{(t)}),\phi_{i}(w_{2}^{(t)}),\dots,\phi_{i}(w_{n}^{(t)}))\in[\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}]$. A full discussion on the optimization problem (19) can be found in Appendix H.
@@
The problem in (19) serves purely as a theoretical intermediate step for establishing the regret upper bound in Theorem 5.4. In practice, we do not compute $\hat{\psi}_{i,\bm{\theta}}$ for every $\bm{\theta}$. As discussed in Algorithm 1, we only solve (19) once, with $\bm{\theta}=\tilde{\bm{\theta}}_{i}$, to obtain $\hat{\psi}_{i,\tilde{\bm{\theta}}i}$. Moreover, by point (4) in Proposition D.4, the function $\psi_{i,\bm{\theta}}$ is increasing (for $\bm{\theta}$ sufficiently close to $\bm{\theta}_{i}$). Since $h_{s_{i}}$ is also increasing, it follows that $\phi_{i,\bm{\theta}}$ inherits this monotonicity. Nonetheless, we did not include this monotonicity constraint in the definition of $\mathcal{S}_{i}$ in (19). This omission is justified because the theoretical convergence rate in Theorem D.6, and consequently the regret upper bound in Theorem 5.4, is driven entirely by the concavity constraint, which imposes a higher-order smoothness condition than monotonicity. However, we do enforce monotonicity in the algorithmic implementation in Algorithm 1.
@@
Let Assumptions 3.1, 3.4 and 5.1 hold. Then, by Proposition 3.5 and Proposition D.4, $\psi_{i,\bm{\theta}}$ is $s_{i}$-concave, where $s_{i}=c_{i}-1>-1$. Let $m=\pi n$, where $\pi=\int_{\mathcal{U}}f_{w}(u)du$ is the proportion of data points $\{w_{i}^{(t)}\}_{t\in[n]}$ lying in the set $\mathcal{U}$ as $n\rightarrow\infty$. For every $\gamma>4$ there exists $m_{0}$, a constant $\mathscr{C}_{i}>0$ and $\widetilde{\mathscr{C}}_{i}>0$ depending only on constants in the assumptions such that
@@
Proof of Theorem D.6.
We first prove the concentration inequality in (20).
@@
holds with probability at least $1-1/2(n+2)^{\gamma}$ for some $\gamma>0$, provided that $n$ is sufficiently large and $\epsilon_{n}\to 0$ as $n\to\infty$. Now, we apply Theorem H.5, using Remark H.6, with the following substitutions:
@@
where $h_{s_{i}}$ is defined in (4). Intersecting $\mathscr{Q}$ with the high-probability event in Theorem H.5 yields the desired result. The only remaining step is to verify that the assumptions of Theorem D.6 imply those of Theorem H.5.
@@
    $\psi_{i,\bm{\theta}}$ is $s_{i}$-concave for some $s_{i}>-1$. According to Proposition D.4, this condition is satisfied provided $\psi_{i}$ is $s_{i}$-concave for some $s_{i}>-1$, which holds by Proposition 3.5 and 3.4, with $s_{i}=c_{i}-1$.
@@
    H.2 holds uniformly in $\bm{\theta}\in\mathbb{S}_{N-1}$ because the design points are drawn from an i.i.d. elliptically symmetric distribution with density generator $g>0$ and then $f_{\mathbf{p}}>0$ on any bounded interval. Thus, $\inf_{\bm{\theta}\in\mathbb{S}_{N-1}}\inf_{u\in\mathcal{U}}f_{\bm{\theta}^{\top}\mathbf{p}}(u)\geq C_{\mathscr{D}}$ for some $C_{\mathscr{D}}>0$.
@@
    H.3 holds because $y_{i}^{(t)}$ are sub-gaussian (indeed $\psi_{i}$ is bounded and $\varepsilon_{i}^{(t)}$ are sub-gaussian) and $\psi_{i,\bm{\theta}}$ is bounded uniformly in $\bm{\theta}\in\mathbb{S}_{N-1}$ by Proposition D.4.
@@
    We prove that H.4 holds uniformly in $\bm{\theta}$. Recall that $\phi_{i,\bm{\theta}}=d_{s_{i}}\circ\psi_{i,\bm{\theta}}$ (where $d_{s_{i}}\triangleq h^{-1}_{s_{i}}$ is defined in (4)). Then $\psi_{i,\bm{\theta}}^{\prime}(u)=d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(u))\psi^{\prime}_{i,\bm{\theta}}(u)$. For every $u,v\in\mathcal{U}$ and $\bm{\theta}\in\mathbb{S}_{N-1}$ we have
@@
    The first component $|d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(u))-d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(v))|\leq L_{d^{\prime}}|u-v|$ by Section H.1, where $L_{d^{\prime}}=\sup_{x}d^{\prime\prime}_{s_{i}}(x)$ depends only on $s_{i}$ and bound of $\psi_{i}$. The component $|\psi^{\prime}_{i,\bm{\theta}}(u)|$ is bounded uniformly by Proposition D.4 point (4). The other component $|\psi^{\prime}_{i,\bm{\theta}}(u)-\psi^{\prime}_{i,\bm{\theta}}(v)|\leq L|u-v|$ for some $L>0$ independent of $\bm{\theta}$ by Proposition D.4 point (4) and $|d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(v))|$ is uniformly bounded by Section H.1. This proves that $\psi_{i,\bm{\theta}}^{\prime}$ is Lipschitz uniformly in $\bm{\theta}\in\mathbb{S}_{N-1}$.
To prove (22) we apply the triangular inequality
@@
and for the first component we apply the concentration inequality in (20) while for the second we use point (3) of Proposition D.4
@@
and then the concentration inequality in Proposition D.3, while retaining the dominating term $O((\log(m)/m)^{2/5})$. ∎
Appendix E Proof of Proposition D.3
@@
w.p. at least $1-2e^{-c\varsigma_{\min}^{2}n/16}$ (for $n$ sufficiently large) for certain constants $c>0$ that depends only on the variance proxy $\sigma_{x}$ of the sub-Gaussian random variable $\mathbf{x}^{(t)}$, and by Lemma E.2 the RHS satisfies
@@
with probability at least $1-\frac{2N}{n}$ (as long as $n\geq N$), where $\mathscr{B}_{i}^{**}$ is a constant defined in (23) that depends solely on $\lambda_{i}$ and the variance proxies of the sub-Gaussian random variables $y_{i}^{(t)}$ and $\mathbf{x}^{(t)}$. Putting the two inequalities together we get with probability at least $1-2e^{-c\varsigma_{\min}^{2}n/16}-\tfrac{2N}{n}$ that, for $n$ sufficiently large
@@
where $\mathscr{B}_{i}^{*}=2\frac{\mathscr{B}_{i}}{\varsigma_{\min}}$. Now we conclude. Recall that $\lambda_{i}>0$ where $\lambda_{i}$ is defined in (25). Let
@@
is a constant that depends solely on $\lambda_{i}$ (defined in (24)) and the variance proxies $\sigma_{y_{i}}$ and $\sigma_{x}$ of the sub-Gaussian random variables $y_{i}^{(t)}$ and $\mathbf{x}^{(t)}$, respectively.
Proof of Lemma E.2.
@@
Appendix F Proof of Proposition D.4
@@ F.1 Proof of (1)
Then $\mathbf{y}^{\top}\mathbf{y}=y_{1}^{2}+\|\bm{\alpha}\|_{2}^{2}$. From the constraint in Equation 27:
@@
where we defined the conditional center (the ”shift”)
@@ F.2 Proof of (2)
where we recall from Equation 31 that $\mathbf{m}_{u}=\mathbf{m}+\Lambda\bm{\theta}\frac{\left(u-\bm{\theta}^{\top}\mathbf{m}\right)}{\bm{\theta}^{\top}\Lambda\bm{\theta}}$. Note that
@@
where we used that $\bm{\theta_{i}},\bm{\theta}\in\mathbb{S}_{N-1}$. By (F.1) we have
@@
proving that $\psi_{i,\bm{\theta}}\in\mathrm{C}^{2}(\mathcal{U})$ for all $\bm{\theta}$. By 3.1 we have that $0<\underline{B}_{\psi_{i}}\leq\psi_{i}\leq\bar{B}_{\psi_{i}}$, which immediately implies that $0<\underline{B}_{\psi_{i}}\leq\psi_{i,\bm{\theta}}\leq\bar{B}_{\psi_{i}}$ uniformly in $\bm{\theta}$, proving point (2).

### v5

@@ Appendix A Summary of individual information
We now construct an estimator of $\psi_{i,\bm{\theta}}$ and establish, in Theorem˜D.6, a uniform convergence result over any bounded subset $K\subset\mathcal{U}$ (where $\mathcal{U}$ is defined in equation 5). Our approach draws on techniques from mosching2020monotone; dumbgen2004consistency, which study uniform convergence in the context of shape-constrained estimation. A key assumption in these works – one that does not require smoothness of the unknown nonparametric function – is that the density of the design points, $f_{w}$, is bounded away from zero on $\mathcal{U}$. This ensures that the design points $w^{(t)}_{i}\in\mathcal{U}$ are asymptotically dense within every subinterval of $\mathcal{U}$, which in turn is sufficient for establishing uniform convergence of the estimator on such intervals. In our setting, $f_{w}$ is supported on all of $\mathbb{R}$, and thus it is bounded away from zero on any bounded set $K$. Consequently, the uniform convergence result applies to any bounded subset $K\subset\mathcal{U}$.
Before establishing the convergence result, we must first characterize the shape constraints satisfied by the family of functions $\{\psi_{i,\bm{\theta}}(\cdot):\bm{\theta}\in\mathbb{S}_{N-1}\}$. In Proposition˜D.4 (with proof deferred to Appendix˜F), we derive a closed-form expression for $\psi_{i,\bm{\theta}}(\cdot)$ and show that both the monotonicity and the $s_{i}=c_{i}-1$-concavity of the original function $\psi_{i}$ are preserved by $\psi_{i,\bm{\theta}}$, uniformly over $\bm{\theta}\in\mathbb{S}_{N-1}$.
@@
Under ˜3.1 and ˜5.1 we have the following properties:
@@
    If also ˜3.4 holds (i.e. $\psi_{i}$ is $s_{i}$-concave for some $s_{i}>-1$ as it arises from Proposition˜3.5) then $\psi_{i,\bm{\theta}}$ is $s_{i}$-concave for any $\bm{\theta}\in\mathbb{S}_{N-1}$.
By definition of $s$-concavity, from point (5) in Proposition˜D.4 we have that for every $\bm{\theta}\in\mathbb{S}_{N-1}$ there exists a concave function $\phi_{i,\bm{\theta}}$ that $\psi_{i,\bm{\theta}}=h_{s_{i}}\circ\phi_{i,\bm{\theta}}$, where $h_{s_{i}}$ is defined in equation 4, and $s_{i}=c_{i}-1$ is known by ˜3.4. We could then estimate $\phi_{i,\bm{\theta}}(u)$ using LSE under concavity restriction, that is
@@
where we recall $w^{(t)}_{i}\triangleq\langle\bm{\theta},\mathbf{p}^{(t)}\rangle$. The interval $\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}]$ is such that $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}]=h_{s_{i}}([\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}])$. Note that $\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}]$ is known because $\underline{B}_{\psi_{i}}$ and $\bar{B}_{\psi_{i}}$ are known by ˜3.1 (recll that, by point (2) in Proposition˜D.4, the bounds $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}$ are inherited from $\psi_{i}$ to $\psi_{i,\bm{\theta}}$). We want to highlight that for any fixed $\bm{\theta}\in\mathbb{S}_{N-1}$, the problem (19) is not infinite dimensional, indeed we only need to recover $n$ variables $(\phi_{i}(w_{1}^{(t)}),\phi_{i}(w_{2}^{(t)}),\dots,\phi_{i}(w_{n}^{(t)}))\in[\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}]$. A full discussion on the optimization problem (19) can be found in Appendix˜H.
@@
The problem in (19) serves purely as a theoretical intermediate step for establishing the regret upper bound in Theorem˜5.4. In practice, we do not compute $\hat{\psi}_{i,\bm{\theta}}$ for every $\bm{\theta}$. As discussed in Algorithm˜1, we only solve (19) once, with $\bm{\theta}=\tilde{\bm{\theta}}_{i}$, to obtain $\hat{\psi}_{i,\tilde{\bm{\theta}}i}$. Moreover, by point (4) in Proposition˜D.4, the function $\psi_{i,\bm{\theta}}$ is increasing (for $\bm{\theta}$ sufficiently close to $\bm{\theta}_{i}$). Since $h_{s_{i}}$ is also increasing, it follows that $\phi_{i,\bm{\theta}}$ inherits this monotonicity. Nonetheless, we did not include this monotonicity constraint in the definition of $\mathcal{S}_{i}$ in (19). This omission is justified because the theoretical convergence rate in Theorem˜D.6, and consequently the regret upper bound in Theorem˜5.4, is driven entirely by the concavity constraint, which imposes a higher-order smoothness condition than monotonicity. However, we do enforce monotonicity in the algorithmic implementation in Algorithm˜1.
@@
Let Assumptions 3.1, 3.4 and 5.1 hold. Then, by Proposition˜3.5 and Proposition˜D.4, $\psi_{i,\bm{\theta}}$ is $s_{i}$-concave, where $s_{i}=c_{i}-1>-1$. Let $m=\pi n$, where $\pi=\int_{\mathcal{U}}f_{w}(u)du$ is the proportion of data points $\{w_{i}^{(t)}\}_{t\in[n]}$ lying in the set $\mathcal{U}$ as $n\rightarrow\infty$. For every $\gamma>4$ there exists $m_{0}$, a constant $\mathscr{C}_{i}>0$ and $\widetilde{\mathscr{C}}_{i}>0$ depending only on constants in the assumptions such that
@@
Proof of Theorem˜D.6.
We first prove the concentration inequality in equation 20.
@@
holds with probability at least $1-1/2(n+2)^{\gamma}$ for some $\gamma>0$, provided that $n$ is sufficiently large and $\epsilon_{n}\to 0$ as $n\to\infty$. Now, we apply Theorem˜H.5, using Remark˜H.6, with the following substitutions:
@@
where $h_{s_{i}}$ is defined in equation 4. Intersecting $\mathscr{Q}$ with the high-probability event in Theorem˜H.5 yields the desired result. The only remaining step is to verify that the assumptions of Theorem˜D.6 imply those of Theorem˜H.5.
@@
    $\psi_{i,\bm{\theta}}$ is $s_{i}$-concave for some $s_{i}>-1$. According to Proposition˜D.4, this condition is satisfied provided $\psi_{i}$ is $s_{i}$-concave for some $s_{i}>-1$, which holds by Proposition˜3.5 and ˜3.4, with $s_{i}=c_{i}-1$.
@@
    ˜H.2 holds uniformly in $\bm{\theta}\in\mathbb{S}_{N-1}$ because the design points are drawn from an i.i.d. elliptically symmetric distribution with density generator $g>0$ and then $f_{\mathbf{p}}>0$ on any bounded interval. Thus, $\inf_{\bm{\theta}\in\mathbb{S}_{N-1}}\inf_{u\in\mathcal{U}}f_{\bm{\theta}^{\top}\mathbf{p}}(u)\geq C_{\mathscr{D}}$ for some $C_{\mathscr{D}}>0$.
@@
    ˜H.3 holds because $y_{i}^{(t)}$ are sub-gaussian (indeed $\psi_{i}$ is bounded and $\varepsilon_{i}^{(t)}$ are sub-gaussian) and $\psi_{i,\bm{\theta}}$ is bounded uniformly in $\bm{\theta}\in\mathbb{S}_{N-1}$ by Proposition˜D.4.
@@
    We prove that ˜H.4 holds uniformly in $\bm{\theta}$. Recall that $\phi_{i,\bm{\theta}}=d_{s_{i}}\circ\psi_{i,\bm{\theta}}$ (where $d_{s_{i}}\triangleq h^{-1}_{s_{i}}$ is defined in equation 4). Then $\psi_{i,\bm{\theta}}^{\prime}(u)=d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(u))\psi^{\prime}_{i,\bm{\theta}}(u)$. For every $u,v\in\mathcal{U}$ and $\bm{\theta}\in\mathbb{S}_{N-1}$ we have
@@
    The first component $|d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(u))-d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(v))|\leq L_{d^{\prime}}|u-v|$ by Section˜H.1, where $L_{d^{\prime}}=\sup_{x}d^{\prime\prime}_{s_{i}}(x)$ depends only on $s_{i}$ and bound of $\psi_{i}$. The component $|\psi^{\prime}_{i,\bm{\theta}}(u)|$ is bounded uniformly by Proposition˜D.4 point (4). The other component $|\psi^{\prime}_{i,\bm{\theta}}(u)-\psi^{\prime}_{i,\bm{\theta}}(v)|\leq L|u-v|$ for some $L>0$ independent of $\bm{\theta}$ by Proposition˜D.4 point (4) and $|d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(v))|$ is uniformly bounded by Section˜H.1. This proves that $\psi_{i,\bm{\theta}}^{\prime}$ is Lipschitz uniformly in $\bm{\theta}\in\mathbb{S}_{N-1}$.
To prove equation 22 we apply the triangular inequality
@@
and for the first component we apply the concentration inequality in equation 20 while for the second we use point (3) of Proposition˜D.4
@@
and then the concentration inequality in Proposition˜D.3, while retaining the dominating term $O((\log(m)/m)^{2/5})$. ∎
Appendix E Proof of Proposition˜D.3
@@
w.p. at least $1-2e^{-c\varsigma_{\min}^{2}n/16}$ (for $n$ sufficiently large) for certain constants $c>0$ that depends only on the variance proxy $\sigma_{x}$ of the sub-Gaussian random variable $\mathbf{x}^{(t)}$, and by Lemma˜E.2 the RHS satisfies
@@
with probability at least $1-\frac{2N}{n}$ (as long as $n\geq N$), where $\mathscr{B}_{i}^{**}$ is a constant defined in equation 23 that depends solely on $\lambda_{i}$ and the variance proxies of the sub-Gaussian random variables $y_{i}^{(t)}$ and $\mathbf{x}^{(t)}$. Putting the two inequalities together we get with probability at least $1-2e^{-c\varsigma_{\min}^{2}n/16}-\tfrac{2N}{n}$ that, for $n$ sufficiently large
@@
where $\mathscr{B}_{i}^{*}=2\frac{\mathscr{B}_{i}}{\varsigma_{\min}}$. Now we conclude. Recall that $\lambda_{i}>0$ where $\lambda_{i}$ is defined in equation 25. Let
@@
is a constant that depends solely on $\lambda_{i}$ (defined in equation 24) and the variance proxies $\sigma_{y_{i}}$ and $\sigma_{x}$ of the sub-Gaussian random variables $y_{i}^{(t)}$ and $\mathbf{x}^{(t)}$, respectively.
Proof of Lemma˜E.2.
@@
Appendix F Proof of Proposition˜D.4
@@ F.1 Proof of (1)
Then $\mathbf{y}^{\top}\mathbf{y}=y_{1}^{2}+\|\bm{\alpha}\|_{2}^{2}$. From the constraint in Equation˜27:
@@
where we defined the conditional center (the "shift")
@@ F.2 Proof of (2)
where we recall from Equation˜31 that $\mathbf{m}_{u}=\mathbf{m}+\Lambda\bm{\theta}\frac{\left(u-\bm{\theta}^{\top}\mathbf{m}\right)}{\bm{\theta}^{\top}\Lambda\bm{\theta}}$. Note that
@@
where we used that $\bm{\theta_{i}},\bm{\theta}\in\mathbb{S}_{N-1}$. By equation F.1 we have
@@
proving that $\psi_{i,\bm{\theta}}\in\mathrm{C}^{2}(\mathcal{U})$ for all $\bm{\theta}$. By ˜3.1 we have that $0<\underline{B}_{\psi_{i}}\leq\psi_{i}\leq\bar{B}_{\psi_{i}}$, which immediately implies that $0<\underline{B}_{\psi_{i}}\leq\psi_{i,\bm{\theta}}\leq\bar{B}_{\psi_{i}}$ uniformly in $\bm{\theta}$, proving point (2).